the owner of candy store mixed some peanuts worth $3per pound some cashews worth $9 per pound and some brazil nuts worth $9 per pound to get 50 pounds of a mixture that would sell for $6 per pounds she used 15 fewer pounds of cashews than peanuts how many pounds of each did she use?
Answer by QH(50)

First assign variables for each nut. Since it is only asking for the pounds of each, the costs are irrelevant.
Peanuts= x
Cashews= x-15
Brazil Nuts= 50-(x+x-15) or 2x-35
Add all of the "nuts" together to equal 50lbs
(x)+(x-15)+(2x-35)=50
4x-50=50
4x=100
x=100/4
(peanuts) x=25 lbs
(cashews) x-15=10 lbs
Brazils) 2x-35=15 lbs
check: 20+10+15=50
50=50